OpenStudy (music0203):

suppose that the amount of algae in a pound doubles every 6 hours. If the pound initially contains 50 pounds of algae, how much algae will be in the pond after 18 hours?

OpenStudy (music0203):

please help me! :)

OpenStudy (music0203):

I will give medals, if someone helps and tells me how

OpenStudy (music0203):

@Michele_Laino

OpenStudy (music0203):

@jim_thompson5910

OpenStudy (michele_laino):

hint: we start with 50 pounds, then after 6 hour we got: 50*2=100 pounds after 12 hours we got: 100*2=200 pounds, and after 18 hours, we got: 200*2=...?

OpenStudy (music0203):

400? :)

OpenStudy (michele_laino):

yes!

OpenStudy (music0203):

thank you so much! how do i give medals??

OpenStudy (michele_laino):

you have to click on the "Best Response" button

OpenStudy (michele_laino):

thanks! :)